OpenStudy (anonymous):

help. ! Write a second equation for the system so that the system will have no solution. 3x+2y=-8

OpenStudy (anonymous):

If \(3x+2y=-8\) then 3x+2y can't also add up to something other than -8.

OpenStudy (anonymous):

it says write a short answer. what should I put. no solution.

OpenStudy (anonymous):

I'm telling you how to figure out the answer.

OpenStudy (anonymous):

okay I don't understand what you wrote as you can see Im not good in math. sorry

OpenStudy (anonymous):

What I'm saying is that if \(3x + 2y = -8\) then -8 is the only thing that 3x+2y can equal. If the system was \[3x + 2y = -8\]\[3x + 2y = 12\] then you'd be saying that 3x+2y simultaneously equals two different things. This is impossible.. thus giving the system no solution.

OpenStudy (anonymous):

okay thank you ! I understand